# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

### v4

@@ III-D Inter-Modality Feature Enhancement and Scaling
where $E^{I}$ and $E^{C}$ are the resulting scaled feature matrices for the respective modalities, enhancing the representations learned by the GAT encoders, and $\odot$ denotes element-wise multiplication. This operation mitigates the effects of disparities in scale or distribution shape between the two sets of features.
III-E Cross-Graph Modal Fusion
To integrate features from cross-graph modality representations at varying scales, we concatenate each image embedding $E^{I}$ with its corresponding representation $\mathcal{H}^{I}$ to form $\hat{Z}^{I}=[\mathcal{H}^{I}\mathbin{\|}E^{I}]$, and similarly for clinical data, $\hat{Z}^{C}=[\mathcal{H}^{C}\mathbin{\|}E^{C}]$, resulting in final embedding matrices. We then combine these cross-modal embeddings through element-wise addition and apply a sigmoid function to create a common space: $\hat{Z}=\sigma(\hat{Z}^{I}+\hat{Z}^{C})$. This process aligns representations from both modalities in a common feature space. Furthermore, to integrate the feature spaces of images and meta-features in the same embedded space, we construct a similarity matrix $\textbf{S}\in\mathbb{R}^{N\times N}$ for each pair of similar patients using the final embedding $\hat{Z}$ learned from the model. The similarity between the $i$-th and $j$-th patients can be defined as follows:
$S_{ij}=\hat{Z}_{i}\cdot(\hat{Z}_{j})^{T},\forall i,j\in\left[1,N\right],$ (13)
where $S_{ij}$ represents the similarity between patients $i$ and $j$, incorporating information from both modalities.
III-F Contrastive Loss Optimization
[…]
@@ III-F Contrastive Loss Optimization
where $y_{i}$ is the one-hot vector of the true label, and $\hat{y}_{i}^{I}$ and $\hat{y}_{i}^{C}$ are the model’s outputs for the image and meta-features, respectively. During the optimization process, we developed a comprehensive loss function that integrates cross-entropy and contrastive loss from the two cross-graph modalities. To further enhance the effectiveness of this combined loss, we incorporated the mean squared error between the similarity matrix $S$ and the diagonal matrix $D_{ii}=\sum_{i}A_{ii}$ when calculating the clustering of the module output fusion. The extend diagonal loss is expressed as:
@@
We use $\beta$ as a leverage coefficient to control the optimization weight of the overall loss, which is defined as follows:
$\mathcal{L}_{CGMCL}=(1-\beta)(\mathcal{L}_{m}+\mathcal{L}_{f})+\beta\mathcal{L}_{contrastive}+\mathcal{L}_{diag},$ (22)
where $\beta$ can be set between 0 and 1, and it controls the contribution level of different losses. The coefficient $\beta$ adjusts the weight assigned to each loss component.
TABLE I: Summary statistics of demographics in PD patients with different subtypes
@@
IV Dataset Collection
@@
  • Parkinson’s disease (PD) Data for this study was collected at Kaohsiung Chang Gung Memorial Hospital, Taiwan, from January 2017 to June 2019, involving 416 patients [27]. The study received approval from the Chang Gung Medical Foundation Institutional Review Board, and all data were de-identified. Four expert nuclear medicine physician annotated the data, labeling PD across three subtypes: Normal, Mildly Abnormal (MA), and Abnormal. Details regarding the annotation criteria can be found in Supplementary Fig. S2, while Table I summarizes the descriptive statistics. The images used in this study were Tc99m TRODAT single-photon emission computed tomography (SPECT) scans acquired using a hybrid SPECT/CT system (Symbia T, Siemens Medical Solution). Image acquisition involved 30-second steps across 120 projections, covering a full 360-degree circular rotation with low-energy, high-resolution parallel-hole collimators. After reconstruction, CT-based attenuation correction imported the images into DaTQUANT for automatic semi-quantification of the DaTQUANT meta-features [28]. Twelve parameters were obtained from DaTQUANT: Striatum Right (S-R), Striatum Left (S-L), Anterior Putamen Right (AP-R), Anterior Putamen Left (AP-L), Posterior Putamen Right (PP-R), Posterior Putamen Left (PP-L), Caudate Right (C-R), Caudate Left (C-L), Putamen/Caudate Ratio Right (P/C-R), Putamen/Caudate Ratio Left (P/C-L), Putamen Asymmetry (PA), and Caudate Asymmetry (CA). Supplementary Fig. S1 contains additional details about these meta-features. The original SPECT images (800 × 1132) were resized to a standardized resolution of 128 × 128 pixels for model development. A total of 412 preprocessed images and their twelve associated quantitative DaTQUANT meta-features were utilized for model training (n = 300) and testing (n = 112).
@@
  • Melanoma dataset [29] The melanoma open dataset utilized for this study is a publicly available 7-point multimodal dataset comprising dermoscopic images and clinical data from 413 training and 395 testing samples. The classification task involves seven key image-based features: 1) pigment network (PN), 2) blue whitish veil (BWV), 3) vascular structures (VS), 4) pigmentation (PIG), 5) streaks (STR), 6) dots and globules (DaG), and 7) regression structures (RS). Additionally, the dataset includes five diagnostic categories: 1) basal cell carcinoma (BCC), 2) blue nevus (NEV), 3) melanoma (MEL), 4) miscellaneous (MISC), and 5) seborrheic keratosis (SK). The dermoscopic images have a resolution of 512 × 768 pixels, while the clinical data contains information on the patient’s gender and lesion location. More melanoma categories and their annotation information can be referenced in Table 1 of the literature  [29]
V EXPERIMENTS
V-A Baseline Methods Comparison
To fairly compare the effectiveness of CGMCL across various methods for both the PD and melanoma datasets, we conducted a quantitative analysis of baseline method comparisons. The results for PD, segmented into three subtypes, are presented in Table II through Table IV, and for the melanoma multi-class dataset in Table V. Initially, we employed conventional machine learning algorithms to assess the impact of a single modality on meta-features. This analysis applied logistic regression, XGBoost, random forest, support vector machines (SVM) [30], and AdaBoost [31] to both datasets. To further evaluate the performance of our proposed CGMCL model in PD classification, we compared it to unimodal two-layer CNN models. We then extended the analysis using SPECT images with three slices to compare different 3D-CNN architectures. When comparing machine learning methods that utilize only meta-features against CNN-based unimodal approaches, nonlinear logistic and tree-based methods (i.e., XGBoost and AdaBoost) achieved accuracies between 0.58 and 0.63 for distinguishing Normal from MA in PD. This indicates that relying solely on quantitative data may result in losing other informative features, such as those derived from images. In contrast, unimodal image feature extraction models substantially improved classification accuracy for the three PD subtypes. Specifically, ResNet18 increased accuracy by 16% for Normal vs. MA, 12.5% for Normal vs. Abnormal, and 1.16% for Normal vs. Abnormal compared to logistic regression. Conversely, for the melanoma dataset, meta-features in ML methods outperform in the DaG, PN, and VS categories.
V-B Compare with the SOTA Methods
Compared with state-of-the-art (SOTA) methods, we examined the performance of the MHCA, DeAF, and TriFormer models, which employ 3D tensor-based CNNs as feature extractors combined with meta-features for multimodal training. Our experimental results yielded the following findings:
  1. 1.
  2. 2.
    We further evaluated the cross-modal graph fusion capability of our proposed CGMCL model in classifying the three PD subtypes. From Table II, using GAT as the cross-modal fusion method with ResNet18 achieves metric values between 0.73 and 0.76, comparable to the SOTA model using 3D-CNNs or unimodal approaches (ACC $\uparrow$ 4.1%, PPV $\uparrow$ 3.9%). This indicates that the fusion of clinical and imaging data can more effectively assist in identifying difficult-to-diagnose early PD symptoms. Moreover, there is a noticeable improvement in classification performance for the other two subtypes (e.g., MA vs. Abnormal and Normal vs. Abnormal).
  3. 3.
@@ V-B Compare with the SOTA Methods
TABLE II: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs MA. The best performance results are highlighted in bold, and “_” indicates the second-best methods
@@
MHCA [32] - 3D-CNN 0.58 $\pm$ 0.02 0.52 $\pm$ 0.02 0.68 $\pm$ 0.03 0.68 $\pm$ 0.03 0.51 $\pm$ 0.02
DeAF [33] - 3D-CNN 0.62 $\pm$ 0.02 0.51 $\pm$ 0.11 0.76 $\pm$ 0.10 0.76 $\pm$ 0.06 0.54 $\pm$ 0.03
TriFormer [12] Transformer 3D-CNN 0.63 $\pm$ 0.02 0.56 $\pm$ 0.04 0.73 $\pm$ 0.02 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02
@@
TABLE III: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Abnormal vs Abnormal. The best performance results are highlighted in bold, and “_” indicates the second-best methods
@@
MHCA [32] - 3D-CNN 0.80 $\pm$ 0.03 0.74 $\pm$ 0.04 0.88 $\pm$ 0.04 0.90 $\pm$ 0.03 0.71 $\pm$ 0.04
DeAF [33] - 3D-CNN 0.87 $\pm$ 0.03 0.84 $\pm$ 0.06 0.91 $\pm$ 0.03 0.93 $\pm$ 0.02 0.80 $\pm$ 0.06
TriFormer [12] Transformer 3D-CNN 0.83 $\pm$ 0.03 0.79 $\pm$ 0.03 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.75 $\pm$ 0.03
@@
TABLE IV: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs Abnormal. The best performance results are highlighted in bold, and “_” indicates the second-best methods
@@
MHCA [32] - 3D-CNN 0.87 $\pm$ 0.01 0.92 $\pm$ 0.03 0.82 $\pm$ 0.02 0.85 $\pm$ 0.01 0.90 $\pm$ 0.03
DeAF [33] - 3D-CNN 0.88 $\pm$ 0.01 0.87 $\pm$ 0.02 0.89 $\pm$ 0.01 0.90 $\pm$ 0.00 0.86 $\pm$ 0.02
TriFormer [12] Transformer 3D-CNN 0.83 $\pm$ 0.03 0.86 $\pm$ 0.03 0.80 $\pm$ 0.03 0.83 $\pm$ 0.03 0.84 $\pm$ 0.03
@@
TABLE V: Comparison of the accuracy of multi-class classification performance between unimodal and proposed CGMCL on melanoma dataset. The best performance results are highlighted in bold, and “_” indicates the second-best methods
@@
ResNet18 CNN 0.84$\pm$0.01 0.48$\pm$0.14 0.59$\pm$0.11 0.65$\pm$0.09 0.74$\pm$0.01 0.64$\pm$0.09 0.82$\pm$0.03 0.75$\pm$0.20
2-layer CNN 0.80$\pm$0.01 0.43$\pm$0.12 0.60$\pm$0.12 0.56$\pm$0.08 0.71$\pm$0.01 0.69$\pm$0.1 0.80$\pm$0.03 0.72$\pm$0.20
7-point [29] CNN 0.85 0.60 0.63 0.69 0.77 0.74 0.82 0.73
HcCNN [34] CNN 0.87 0.66 0.69 0.71 0.81 0.72 0.85 0.74
AMFAM [35] GAN 0.88 0.64 0.71 0.71 0.81 0.75 0.83 0.75
FusionM4Net [36] CNN 0.89$\pm$0.00 0.66$\pm$0.02 0.72$\pm$0.01 0.69$\pm$0.01 0.81$\pm$0.01 0.76$\pm$0.01 0.82$\pm$0.01 0.76$\pm$0.01

### v5

@@ 3.4 Inter-Modality Feature Enhancement and Scaling
where $E^{I}$ and $E^{C}$ are the resulting scaled feature matrices for the respective modalities, enhancing the representations learned by the GAT encoders, and $\odot$ denotes element-wise multiplication. Intuitively, IMFES can adaptively reweight to balance modality contributions and mitigate dominance from any single source.
3.5 Cross-Graph Modal Fusion
After feature refinement, the embeddings from both graphs are projected into a shared latent space through element-wise fusion. Specifically, each image-graph embedding $E^{I}$ and its corresponding graph representation $\mathcal{H}^{I}$ are first combined to form $\hat{Z}^{I}=[\mathcal{H}^{I}\mathbin{\|}E^{I}]$, while the clinical-graph embeddings $E^{C}$ and $\mathcal{H}^{C}$ are combined in the same way as $\hat{Z}^{C}=[\mathcal{H}^{C}\mathbin{\|}E^{C}]$. Then, the final image-graph embedding $\hat{Z}^{I}$ and the clinical-graph embedding $\hat{Z}^{C}$ are aligned by combining them via element-wise interaction and a nonlinear activation $\sigma(\cdot)$ to construct a unified latent space: $Z=\sigma(\hat{Z}^{I}+\hat{Z}^{C}).$ To further capture inter-patient relationships within this unified latent space, a similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ is constructed to quantify pairwise correlations between patients:
$S_{ij}=Z_{i}\cdot(Z_{j})^{T},\forall i,j\in\left[1,N\right],$ (13)
where $S_{ij}$ encodes the joint similarity derived from both imaging and clinical modalities.
3.6 Contrastive Loss Optimization
To enforce consistency across modalities, CGMCL introduces a cross-graph contrastive loss that simultaneously maximizes the similarity between positive patient pairs and increases the separation between dissimilar ones. The positive mask matrix is defined as $D_{pos}=\Theta(\hat{A}^{I}+\hat{A}^{C})$. In contrast, the negative mask matrix is defined as $D_{neg}=\Theta\left((1-\hat{A}^{I})+(1-\hat{A}^{C})\right)$, where $\hat{A}^{I}$ and $\hat{A}^{C}$ are adjacency matrices with self-loops. Here, $\Theta(\cdot)$ represents a threshold function, which can be expressed as:
@@ 3.6 Contrastive Loss Optimization
For each modality processed within the graph-modal encoder, the modality-specific feature representations are independently projected into the label space via dedicated MLP classification heads. Formally, this process can be expressed as
$\hat{y}_{i}^{I}=\text{MLP}_{I}(\hat{Z}^{I}_{i}),$ (19)
$\hat{y}_{i}^{C}=\text{MLP}_{C}(\hat{Z}^{C}_{i}),$ (20)
where $\hat{y}_{i}^{I}$ and $\hat{y}_{i}^{C}$ denote the predicted class probability vectors produced by the image–graph encoder and clinical–graph encoder, respectively. Subsequently, for the final multi-modality prediction, the representations from each modality are combined and passed through a softmax activation as follows:
$\hat{y}^{*}=\arg\max\big(\text{softmax}(\hat{Z}^{I}+\hat{Z}^{C})\big).$ (21)
Here, $\hat{y}^{*}$ represents the final predicted class obtained by fusing complementary information from the image and clinical domains through a soft voting mechanism, where combined confidence scores from both modalities jointly determine the final decision and improve robustness.
@@
where $y_{i}$ is the one-hot vector of the true label, and $\hat{y}_{i}^{I}$ and $\hat{y}_{i}^{C}$ are the model’s outputs for the image and meta-features, respectively.
During the optimization process, we developed a comprehensive loss function that integrates cross-entropy and contrastive loss from the two cross-graph modalities. To further enhance the effectiveness of this combined loss, we incorporated the mean squared error between the similarity matrix $S$ and the diagonal matrix $D_{ii}=\sum_{i}A_{ii}$ when calculating the clustering of the module output fusion. The extend diagonal loss is expressed as:
@@
where $\mathcal{L}_{diag}$ enforces structural consistency between graphs. We use $\beta$ as a leverage coefficient to control the optimization weight of the overall loss, which is defined as follows:
$\mathcal{L}_{CGMCL}=(1-\beta)(\mathcal{L}_{I}+\mathcal{L}_{C})+\beta\mathcal{L}_{contrastive}+\mathcal{L}_{diag},$ (25)
where $\beta$ can be set between 0 and 1, and it controls the contribution level of different losses. The coefficient $\beta$ adjusts the weight assigned to each loss component. Specifically, $\mathcal{L}_{I}$ and $\mathcal{L}_{C}$ represent the supervised cross-entropy classification losses for the image modality and clinical meta-feature modality, respectively, as defined in Eq. 22 and Eq. 23.
Table 1: Summary statistics of demographics in PD patients with different subtypes
@@
Figure 4: Three subtypes of PD annotation. Early abnormalities typically manifest as unilateral putamen decline affecting the P/C ratio and symmetry, with progression involving AP and caudate until bilateral symmetric reduction; overall status is summarized by S = C + AP + $P$.
4 Dataset Collection
  • 1.
    Parkinson’s disease (PD) Data for this study was collected at Kaohsiung Chang Gung Memorial Hospital, Taiwan, from January 2017 to June 2019, involving 416 patients Ding_2021. The study received approval from the Institutional Review Board, and all data were de-identified. Four expert nuclear medicine physicians annotated the data, labeling PD across three subtypes: Normal, Mildly Abnormal (MA), and Abnormal, as shown in Fig. 4. Table 1 and 2 summarize the descriptive statistics of the demographics and meta-features.
@@ 4 Dataset Collection
    The images used in this study were Tc99m TRODAT single-photon emission computed tomography (SPECT) scans acquired using a hybrid SPECT/CT system (Symbia T, Siemens Medical Solution). Image acquisition involved 30-second steps across 120 projections, covering a full 360-degree circular rotation with low-energy, high-resolution parallel-hole collimators. After reconstruction, CT-based attenuation correction imported the images into DaTQUANT for automatic semi-quantification of the DaTQUANT meta-features Brogley_2019. Twelve parameters were obtained from DaTQUANT: Striatum Right (S-R), Striatum Left (S-L), Anterior Putamen Right (AP-R), Anterior Putamen Left (AP-L), Posterior Putamen Right (PP-R), Posterior Putamen Left (PP-L), Caudate Right (C-R), Caudate Left (C-L), Putamen/Caudate Ratio Right (P/C-R), Putamen/Caudate Ratio Left (P/C-L), Putamen Asymmetry (PA), and Caudate Asymmetry (CA). The Fig. 1 contains additional details about these meta-features. The original SPECT images (800 × 1132) were resized to a standardized resolution of 128 × 128 pixels for model development. A total of 412 preprocessed images and their twelve associated quantitative DaTQUANT meta-features were utilized for model training (n = 300) and testing (n = 112).
    Table 2: The mean ± standard deviation of twelve meta-features for three PD subtypes across different regions
    Meta-features Normal MA Abnormal
    S-R 0.78 ± 0.22 0.65 ± 0.21 0.35 ± 0.16
    S-L 0.78 ± 0.22 0.65 ± 0.21 0.36 ± 0.15
    AP-R 0.81 ± 0.23 0.65 ± 0.23 0.32 ± 0.16
    AP-L 0.83 ± 0.24 0.66 ± 0.22 0.35 ± 0.15
    PP-R 0.60 ± 0.25 0.43 ± 0.22 0.23 ± 0.15
    PP-L 0.58 ± 0.25 0.42 ± 0.22 0.23 ± 0.12
    C-R 0.82 ± 0.23 0.73 ± 0.24 0.44 ± 0.21
    C-L 0.83 ± 0.23 0.75 ± 0.24 0.46 ± 0.21
    P/C-R 0.97 ± 0.08 0.93 ± 0.09 0.91 ± 0.09
    P/C-L 0.96 ± 0.08 0.92 ± 0.09 0.91 ± 0.09
    PA 0.04 ± 0.03 0.04 ± 0.04 0.06 ± 0.05
    CA 0.04 ± 0.03 0.05 ± 0.04 0.06 ± 0.05
  • 2.
@@
    Melanoma dataset kawahara2018seven The melanoma open dataset utilized for this study is a publicly available 7-point multimodal dataset comprising dermoscopic images and clinical data from 413 training and 395 testing samples. The classification task involves seven key image-based features: 1) pigment network (PN), 2) blue whitish veil (BWV), 3) vascular structures (VS), 4) pigmentation (PIG), 5) streaks (STR), 6) dots and globules (DaG), and 7) regression structures (RS). Additionally, the dataset includes five diagnostic categories: 1) basal cell carcinoma (BCC), 2) blue nevus (NEV), 3) melanoma (MEL), 4) miscellaneous (MISC), and 5) seborrheic keratosis (SK). The dermoscopic images have a resolution of 512 × 768 pixels, while the clinical data contains information on the patient’s gender and lesion location. More melanoma categories and their annotation information can be referenced in Table 1 of the literature  kawahara2018seven
    Model Params(M) GFLOPs Latency(ms) TrainTime(s) GPU(MB)
    ResNet18 11.178 67.880 8.522 31.99 1792.3
    Cluster-GT-T 16.747 68.083 46.526 51.60 1859.5
    CoBFormer 16.892 68.192 9.704 34.49 1960.0
    GTC 18.233 68.952 9.787 34.12 2019.5
    GEAET 18.728 68.315 9.460 33.77 2011.7
    GradFormer 16.508 68.212 9.341 33.57 1970.9
    CGMCL 15.070 ↑ 68.052 ↑ 8.879 ↑ 32.94 ↑ 1997.5 ↓
    Table 3: Performance comparison of models sorted by training time and GPU usage in descending order. Arrows next to CGMCL indicate whether it outperforms (↑) or underperforms (↓) the best non-ResNet model in each metric.
@@
Table 4: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs MA. The best performance results are highlighted in bold, and "_" indicates the second-best methods.
@@
MHCA taylor2019co Transformer 3D-CNN 0.58 $\pm$ 0.02 0.52 $\pm$ 0.02 0.68 $\pm$ 0.03 0.68 $\pm$ 0.03 0.51 $\pm$ 0.02
DeAF li2023deaf Transformer 3D-CNN 0.62 $\pm$ 0.02 0.51 $\pm$ 0.11 0.76 $\pm$ 0.10 0.76 $\pm$ 0.06 0.54 $\pm$ 0.03
TriFormer liu2023triformer Transformer 3D-CNN 0.63 $\pm$ 0.02 0.56 $\pm$ 0.04 0.73 $\pm$ 0.02 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02
Graph-Transformer Models (2-D)
GEAET GEAET_cite Graph-Transformer ResNet18 0.71 $\pm$ 0.03 0.70 $\pm$ 0.06 0.73 $\pm$ 0.06 0.74 $\pm$ 0.05 0.68 $\pm$ 0.05
CoBFormer Xing2024Less Grpah-Transformer 0.72 ± 0.04 0.72 ± 0.06 0.71 ± 0.06 0.73 ± 0.04 0.70 ± 0.05
GTC sun2025gtc Mixture Graph-Transformer 0.73 ± 0.04 0.69 ± 0.07 0.77 ± 0.06 0.78 ± 0.07 0.69 ± 0.05
GradFormer gradformer_cite Graph-Transformer 0.72 ± 0.02 0.69 ± 0.06 0.76 ± 0.05 0.77 ± 0.05 0.69 ± 0.04
Cluster-GT (N2C-Attn-T) huang2024cluster Graph-Transformer 0.72 ± 0.03 0.68 ± 0.07 0.77 ± 0.06 0.77 ± 0.06 0.68 ± 0.05
Cluster-GT (N2C-Attn-L) huang2024cluster Graph-Transformer 0.74 ± 0.02 0.74 ± 0.03 0.75 ± 0.04 0.77 ± 0.04 0.71 ± 0.03
@@
Comparison with Unimodal Improvements (%) - +4.11%∗∗ +0.00% +13.04%∗∗∗ +6.76%∗∗∗ +1.39%
Comparison with Transformer-based Improvements (%) - +20.63%∗∗∗ +33.93%∗∗∗ +2.63% +3.95% +32.73%∗∗∗
Comparison with Graph-Transformer Improvements (%) - +2.70%∗∗ +1.35% +1.30% +1.28% +2.82%
Significance. p << 0.05, ∗∗p << 0.01, ∗∗∗p << 0.001 indicate statistically significant improvements of CGMCL compared to the corresponding baseline models.
Table 5: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Abnormal vs Abnormal. The best performance results are highlighted in bold, and "_" indicates the second-best methods
@@
MHCA taylor2019co Transformer 3D-CNN 0.80 $\pm$ 0.03 0.74 $\pm$ 0.04 0.88 $\pm$ 0.04 0.90 $\pm$ 0.03 0.71 $\pm$ 0.04
DeAF li2023deaf Transformer 3D-CNN 0.87 $\pm$ 0.03 0.84 $\pm$ 0.06 0.91 $\pm$ 0.03 0.93 $\pm$ 0.02 0.80 $\pm$ 0.06
TriFormer liu2023triformer Transformer 3D-CNN 0.83 $\pm$ 0.03 0.79 $\pm$ 0.03 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.75 $\pm$ 0.03
Graph-Transformer Models (2-D)
GEAET GEAET_cite Graph-Transformer ResNet18 0.88 ± 0.02 0.88 ± 0.04 0.88 ± 0.03 0.88 ± 0.03 0.88 ± 0.03
CoBFormer Xing2024Less Graph-Transformer 0.85 ± 0.01 0.84 ± 0.02 0.85 ± 0.03 0.85 ± 0.02 0.84 ± 0.01
GTC sun2025gtc Mixture Graph-Transformer 0.87 ± 0.02 0.86 ± 0.02 0.87 ± 0.02 0.88 ± 0.02 0.86 ± 0.02
GradTransformer gradformer_cite Graph Transformer 0.88 ± 0.01 0.87 ± 0.02 0.89 ± 0.02 0.89 ± 0.02 0.87 ± 0.01
Cluster-GT (N2C-Attn-T) huang2024cluster Graph-Transformer 0.86 ± 0.01 0.84 ± 0.03 0.87 ± 0.04 0.87 ± 0.03 0.85 ± 0.02
Cluster-GT (N2C-Attn-L) huang2024cluster Graph-Transformer - 0.84 ± 0.02 0.85 ± 0.03 0.84 ± 0.01 0.84 ± 0.01 0.85 ± 0.03
@@
Comparison with Unimodal Improvements (%) - +2.30%∗∗ +0.00% +5.81%∗∗∗ +3.41%∗∗ +1.15%
Comparison with Transformer-based Improvements (%) - +2.30%∗∗∗ +4.76%∗∗∗ +0.00% 0.00% +10.00%∗∗∗
Comparison with Graph-Transformer Improvements (%) - +1.14%∗∗ +0.00% +2.25%∗∗ +2.25%∗∗ +0.00%
Significance. p << 0.05, ∗∗p << 0.01, ∗∗∗p << 0.001 indicate statistically significant improvements of CGMCL compared to the corresponding baseline models.
Table 6: The comparison of the proposed CGMCL’s performance (mean $\pm$ std) between unimodal and multimodal approaches for PD subtypes: Normal vs Abnormal. The best performance results are highlighted in bold, and "_" indicates the second-best methods
@@
MHCA taylor2019co Transformer 3D-CNN 0.87 $\pm$ 0.01 0.92 $\pm$ 0.03 0.82 $\pm$ 0.02 0.85 $\pm$ 0.01 0.90 $\pm$ 0.03
DeAF li2023deaf Transformer 3D-CNN 0.88 $\pm$ 0.01 0.87 $\pm$ 0.02 0.89 $\pm$ 0.01 0.90 $\pm$ 0.00 0.86 $\pm$ 0.02
TriFormer liu2023triformer Transformer 3D-CNN 0.83 $\pm$ 0.03 0.86 $\pm$ 0.03 0.80 $\pm$ 0.03 0.83 $\pm$ 0.03 0.84 $\pm$ 0.03
Graph-Transformer Models (2-D)
GEAET GEAET_cite Graph-Transformer ResNet18 0.90 $\pm$ 0.01 0.89 $\pm$ 0.02 0.91 $\pm$ 0.02 0.91 $\pm$ 0.02 0.89 $\pm$ 0.02
CoBFormer Xing2024Less Graph-Transformer 0.90 ± 0.01 0.89 ± 0.01 0.90 ± 0.02 0.90 ± 0.02 0.89 ± 0.01
GTC sun2025gtc Mixture Graph-transformer 0.89 ± 0.01 0.88 ± 0.02 0.90 ± 0.02 0.91 ± 0.02 0.88 ± 0.02
GradFormer gradformer_cite Graph-Transformer 0.88 ± 0.01 0.87 ± 0.02 0.89 ± 0.02 0.89 ± 0.02 0.87 ± 0.02
Cluster-GT-T huang2024cluster Graph-Transformer 0.88 ± 0.01 0.86 ± 0.03 0.91 ± 0.04 0.91 ± 0.03 0.86 ± 0.02
Cluster-GT-L huang2024cluster Graph-Transformer 0.87 ± 0.02 0.86 ± 0.04 0.88 ± 0.02 0.89 ± 0.03 0.86 ± 0.03
@@
Comparison with Unimodal Improvements (%) - +0.00% +1.12% +0.00% +0.00% +0.00%
Comparison with Transformer-based Improvements (%) - +2.27%∗∗∗ +0.00% +2.25%∗∗∗ +1.11%∗∗ +0.00%
Comparison with Graph-Transformer Improvements (%) - +0.00% +1.12% +0.00% +0.00% +0.00%
Significance. p << 0.05, ∗∗p << 0.01, ∗∗∗p << 0.001 indicate statistically significant improvements of CGMCL compared to the corresponding baseline models.
Table 7: Comparison of the accuracy of multi-class classification performance between unimodal and proposed CGMCL on melanoma dataset. The best performance results are highlighted in bold, and "_" indicates the second-best methods
@@
[…]
Multimodal Fusion Methods
7-point kawahara2018seven CNN 0.85 0.60 0.63 0.69 0.77 0.74 0.82 0.73
HcCNN bi2020multi CNN 0.87 0.66 0.69 0.71 0.81 0.72 0.85 0.74
AMFAM wang2022adversarial GAN 0.88 0.64 0.71 0.71 0.81 0.75 0.83 0.75
[…]
Vision-Transformer
SHViT yun2024shvit Vision-Transformer 0.82±0.02 0.56±0.04 0.71±0.05 0.62±0.01 0.70±0.03 0.78±0.03 0.83±0.01 0.93±0.01
[…]